40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 117 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  11067   Wed Feb 25 14:18:28 2015 ranaUpdateLSCDelay line re-installed, measurements round 2

The Schnupp asymmetry is definitely not an important parameter (no need for Koji to explode). It only serves to give us a bigger Q phase signal slope, but is not significant for the I phase signals.

The main anomaly is that the REFL33 optical gain (W/m) seems to have been reduced so much by the phase and amplitude adjustment of the 55 MHz modulation signal. One guess is that the true 3f signal is being made not by the (2*f1 - (-f1)) beat, but by some higher order beat. In addition to the usual RF fields produced by the 2 modulations, we must consider that the sidebands on sidebands produce intermodulation fields just after the EOM and so the fields with which we interrogate the PRMI are more complicated.

Jenne's Optickle calculation today should show us what the sensing matrix contribution is from each pair of fields, so that we can have a sensing matrix signal budget.

 

  11066   Wed Feb 25 12:16:27 2015 KojiUpdateLSCDelay line re-installed, measurements round 2

WHAT? WHAT? WHAT? It's obviously opposite.

If the reflectivity of the front mirror is fixed (=PRM reflectivity), the finesse increases when the reflectivity of the end
mirror (=Compond mirror reflectivity) increases. i.e. 11MHz has higher finesse, 55MHz has lower finesse.

{\cal F} = \frac{\pi \sqrt{r_{\rm PRM} r_{\rm COM}}}{1-r_{\rm PRM} r_{\rm COM}}

If the reflectivity of the front mirror is fixed, the amplitude gain of the cavity is higher when the reflectivity of the end mirror increases. i.e. 11MHz has higher gain, 55MHz has lower gain

g_{\rm PRM} = \frac{t_{\rm PRM}}{1-r_{\rm PRM} r_{\rm COM}}
 

Quote:

Our Schnupp asymmetry is small (3.9cm, IIRC), so the transmission of the 11MHz signal out the dark port is small.  This means that the finesse of the PRC for 11MHz isn't so huge.  On the other hand, since 55MHz is a higher frequency, the transmission out the dark port is larger and is a closer match to the PRM transmission, so the finesse of the PRC for 55MHz is higher. 

 

  11065   Wed Feb 25 11:01:05 2015 manasaUpdateComputer Scripts / ProgramsNew screen for FOL PID loop

Created a new medm screen C1ALS_FOL_PID.adl for FOL PID loop control in /medm/als/master/

This is not currently linked to the sitemap screen.

  11064   Wed Feb 25 04:59:47 2015 JenneUpdateLSCDelay line re-installed, measurements round 2

[Jenne, EricQ, Rana]

We spent this evening measuring and thinking about our 3f signals, and the effect of the modulation cancellation. 

I reinstalled the delay line box, and reduced the modulation depth of the 55MHz signal, so that we are in the state of modulation cancellation - there should be almost no power at 33MHz injected into the vacuum.  I carefully tuned the demod phase of the REFL 11, 33 and 55 MHz PDs, and locked the PRMI on REFL55 I&Q.  The signal in REFL 165 was very tiny, so as best as I could tell, the demod phase that Koji found last week was correct. 

Here is a little record of what the demodulation phases should be, for the "nominal" and "cancellation" configurations, so that we don't have to continually use the time machine.

  "Nominal" configuration 3f modulation cancellation
REFL 11 20 deg 76.0 deg (tuned to nearest 0.5 deg)
REFL 33 142.2 deg 120.3 deg (tuned to nearest 0.1 deg)
REFL 55 19 deg 173.0 deg (tuned to nearest 0.5 deg)
REFL 165 -172 deg 18 deg (same number as last week)
AS 55 -166.1 deg -111.1 deg (same number as last week)

Also, here is the locking recipe for REFL55 I&Q in the cancellation configuration:

PRMI, 3f cancellation, REFL55 I&Q MICH PRCL
Input Matrix acquire with -2*REFL55Q, then put matrix to -15*REFL55Q -12*REFL55I
Gain 3.0 -0.1
DoF trigger POP22I; 50 up, 0.5 down POP22I; 50 up, 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 1, 2, 3, 6, 9; 50 up, 0.5 down, 5 second wait FM 1, 2, 6, 9; 50 up, 0.5 down, 1 second wait
Normalization none none
Output matrix -1*ITMX, +1*ITMY 1*PRM

With this setup, we measured the sensing matrix.  MICH had an excitation at 370.123 Hz with 8,000 counts to -ITMX+ITMY (this is about 0.3nm for each ITM), and PRCL had an excitation at 404.123 Hz with 50 counts to PRM.  For tonight, here is a PDF of the peaks in DTT.  The data is saved in /users/Templates/LSC_error_spectra/SensMat_PRMI_24Feb2015.xml. 

SensMat_PRMI_24Feb2015.pdf


Rana has proposed to us an idea for why the REFL 33 signal should be dominated by the 55*22 contribution, rather than -11*22.  Eric is in the process of checking this out with a Mist model to see if it makes sense. Here's the gist:

Our Schnupp asymmetry is small (3.9cm, IIRC), so the transmission of the 11MHz signal out the dark port is small.  This means that the finesse of the PRC for 11MHz isn't so huge.  On the other hand, since 55MHz is a higher frequency, the transmission out the dark port is larger and is a closer match to the PRM transmission, so the finesse of the PRC for 55MHz is higher. 

Since the magnitudes of the fields at the reflection port are not changing significantly, our PDH signals are being created by the relative phase of something which is anti-resonant (ex. carrier or 22MHz for sideband lock) vs. something which is resonant (ex. 11MHz or 55MHz).  Since the finesse of the 55MHz signal is larger, the accumulated phase change is greater, so we expect a larger slope to our PDH signals that involve 55MHz as compared to those that use 11MHz.

If we are comparing the contributions between -11*22 and 55*22, they both include the anti-resonant 22MHz. So, the difference in the signal strengths comes directly from the difference in phase accumulation due to the variation in the dark port transmission.

EricQ had a thought, and while I have enough awake brain cells to report the thought, we're going to have to revisit it when more of our brains are awake.  In either case, the transmission through the dark port is small compared to the transmission of the ITMs, so why don't the ITMs dominate the finesse calculation, and thus are the 11MHz and 55MHz really getting that much of a difference in finesse?  To be checked out.

Attachment 1: SensMat_PRMI_24Feb2015.pdf
SensMat_PRMI_24Feb2015.pdf
  11063   Wed Feb 25 04:21:58 2015 ericqUpdateCDSSome model updates

I changed the suspension library block to acquire the SUS_[optic]_LSC_OUT channels at 16k for sensing matrix investigations. We could save the FB some load by disabling these and oplev channels in the mode cleaner optic suspensions. 

I removed nonexistant PDs from c1cal, to try and speed it up from its constantly overflowing state. It's still pretty bad, but under 60us most of the time. 

I also cleaned out the unused IPCs for simulated plant stuff from c1scx and c1sus, to get rid of red blinkeys. 

  11062   Tue Feb 24 23:39:16 2015 JenneUpdateLSCDelay line re-installed again

About 5-10 minutes ago I just put in the modulation cancellation setup according to the recipe in http://nodus.ligo.caltech.edu:8080/40m/11032

  11061   Tue Feb 24 18:54:26 2015 ericqUpdateASCSingle arm QPD ASC stability

I've lowered the UGFs for the transmission QPD servos to ~1-2Hz, and made it just an integrator. I left the arms locked with the QPD servos on for a few hours during the daytime today, and they succesfully prevented the Y arm from losing power from alignment drift for ~4 hours. Turning the servo off caused TRY to drop to ~0.6 or so. 

The X arm was only held for 2 hours or so, because after some unlock/drift event the power was below the servo trigger threshold. However, after gently nudging ETMX to get the transmission above the threshold, the servo kicked in, and brought it right back to TRX=1.0

Unfortunately, daqd was dead for much of the day, so I don't have much data to show; the trend was inferred from the wall striptool. 

It is not proven that there aren't further issues that prevent this from working with higher / more dynamic arm powers, but this is at least a point in favor of it working. 

EDIT: Here's a screenshot of the wall StripTool. Brown is TRY, blue is TRX. The downturn at the very end is me deactivating the servos. 

There is no scientific justifcation for the 0.9 threshold. Really, I should look at the noise/SNR again, now that there is some ND filtering on the QPDs. 

Attachment 1: trend.png
trend.png
  11060   Tue Feb 24 01:57:01 2015 JenneUpdateLSCModelled effect of relative modulation phase

I have used Optickle to model the effect of changing the phase between the 11MHz and 55MHz EOMs.  Also, I have looked at what modulation order is most significant (we hope it's -22*11 and -11*22).


First, so that we can compare these numbers more directly to measurements, I have included in the model the fraction of light that gets to each PD.  I'm assuming that the Faraday is about 80% transmissive, but I don't know what the true number is. Here's a cartoon showing the attenuation factors.

EDIT, 26March2015, JCD:  REFL path updated.  See elog 11172.


To model the change in relative phase between the 11MHz and 55MHz modultions, I have held the 11MHz EOM stationary, and moved the 55MHz EOM.  Since I needed an actual distance, I used a conversion factor,

\frac{\rm m}{\rm deg} = \frac{360 \ \rm deg}{55 \ \rm MHz} (\rm speed \ of \ light \ [m/s])

The sensing matrix was measured at 143Hz.  It has been corrected from mevans-meters to Newtons as the denominator.

The big thing to notice here is that the PRCL magnitude is not changing by a factor of 20.  More like a factor of 2.  BUT, I have not yet included the fact that Koji also reduced the 55MHz modulation amplitude by a factor of 3. 

As for the mysterious degeneracy of all the REFL PDs, I think we need to take a more careful measurement.  It's possible that we were seeing the real thing for REFL33, but the others don't seem to change in degeneracy with relative modulation phase. 


Why does it even matter for the 3*f1 signal what is going on with the f2 modulation?  Well, it appears that we are definitely being dominated by the 44*11 and 55*22 components.

To check this, I restricted Optickle to various orders of modulation (ex. up to second order only includes the [-22, -11, 0, 11, 22] MHz components), and plotted them all.  The change in the signal between one trace and another is the contribution from that extra modulation order.  The traces are only minutely different between orders, after the 5th order.  So, since they're all overlapping with the 5th order trace anyway, I don't plot them.

EDIT: to clarify, when I said "up to X order", I meant up to that order in 11MHz sidebands.  Optickle is applying the 11MHz and 55MHz modulations in the same way every time, but then I specify up to what order to include in the summation of different contributions to the field at a given port.  So, for the "up to 2nd order", I only include cross terms that come from [-22, -11, 0, 11, 22] MHz.  For the next order, I only include terms that come from [-33, -22, -11, 0, 11, 22, 33] MHz, etc.  So, there are no 55MHz effects when I'm only including contributions up to 2nd order (since there is a maximum cross beatnote of 44MHz), but starting with 3rd order, I do start to see signals in, for example, REFL55 and AS55, since I get terms from -22*33 and -33*22.  The first order in 55MHz (i.e. 55MHz*Carrier) only starts to show up when I calculate "up to 5th order" and above, since that includes [-55, -44, -33, -22, -11, 0, 11, 22, 33, 44, 55] MHz.


What happens if I reduce the 55MHz modulation depth by 10dB?  Since we are dominated by 55MHz-related signals, the signal at REFL 33 goes down. The maximum change we could have seen for the REFL33 PRCL signal (difference between max of blue trace and min of orange trace) is a factor of 27.


Where are we on the x-axis of these plots, and where was the maximal cancellation place that Koji found?  I need to re-check that part of the code tomorrow, to make sure that I've included all of my contributions from different components of the (field* field) matrix.

But, the moral of the story for tonight is that at least for the REFL33 signal, it's actually plausible that the optical gain went down by a factor of 20, and that the MICH and PRCL signals were degenerate.  I suspect that the total cancellation place that Koji found was somewhere around 175 degrees on the x-axis of these plots and that our nominal place is around 0 deg - around there, both the magnitude and the phase situations are possible simultaneously.

Attachment 1: ChangeInAngle_PRCLvsMICH.png
ChangeInAngle_PRCLvsMICH.png
Attachment 2: REFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.png
REFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.png
Attachment 3: REFL33_SensMat_vsModulationPhase.png
REFL33_SensMat_vsModulationPhase.png
Attachment 4: REFL33_PRCL_nominalVsLoweredGamma.png
REFL33_PRCL_nominalVsLoweredGamma.png
Attachment 5: REFL33_MICH_nominalVsLoweredGamma.png
REFL33_MICH_nominalVsLoweredGamma.png
  11059   Mon Feb 23 21:57:13 2015 KojiUpdateLSCDelay line installed again (experiment, round 1)

Last Wednesday we tried PRMI 3f modulation cancellation. Under the 3f modulation cancellation, the PRMI could not be locked
with REFL signals, and the PRCL signal was just barely sufficient to lock PRCL with help of AS55Q MICH.

- The PRCL signal level in REFL33 was reduced by factor of 20 compared with the conventional modulation setting.
=> The 3f modulation cancellation does not chage the level of 11/22MHz sidebands, it is expected that REFL33 signal
has no significant change of the signal level. But it does.  If we change the relative phase between the modulations
at 11 and 55MHz, the signal level is recovered by factor of 5. Therefore something related to 55MHz modulation
(55MHz x 22MHz, or 44MHz x 11MHz) was contributing more than -11MHzx22MHz.

- Under the 3f demodulation cancellation, MICH signal in the REFL ports were extremely weak and there was
no hope to use it for any feedback control.

- WIth the PRMI locking by REFL33->PRCL and AS55Q->MICH, the sensing matrix was measured. All of the REFL
ports however, showed extremely degenerate sensing matrix between MICH and PRCL.

This was enough confusing to us, and we didn't draw any useful information from these. Here are some ideas to
investigate what is happening in out optical and electrical system.

- One approach is to use as simple optical setup as possible to inspect our sensing systems. For example,
we may want to try PRX/PRY/XARM/YARM cavities to check the functions of the REFL diodes and qualitatively characterize
the sensing chain (Optical gain [W/m], noise level, demodulation phase) so that we can compare these with
an interferometer seinsing model.

- Another approach is to change the mdulation setting more freely and empirically try to find the characteristic
of our optical/electrical systems. e.g. change the relative modulation phase and/or 55MHz attenuation, and try to understand
how 11-22, 11-44, 22-55, 0-33 pairs are contributing the signal.

  11058   Mon Feb 23 15:04:10 2015 ZachUpdateGeneralElog restarted

The elog crashed while I was creating an entry to the Cryo log. I restarted it with the start-elog.csh script.

  11057   Mon Feb 23 08:06:59 2015 ranaUpdateASCQPD frontend code unified

Might have to also get the OL screens that go with this new code to see, but the calibrations don't go into the matrix, but rather into the OLPIT/OLYAW filter banks which follow the division, but before the servo filter banks.

  11056   Fri Feb 20 19:09:48 2015 ericqUpdateASCQPD frontend code unified

I have changed all of the oplevs and transmon QPDs to use the common ISC QPD library block, which differs mainly in its divide by zero protection. 

c1scx.mdl and c1scy.mdl were directly changed for the transmon QPDs. The oplevs were done by changing the sus_single_control.mdl library part, which is used for all of the SOSs. 

Then, because of the underscore introduced (i.e. OLPIT becomes OL_PIT because there is an OL block), I went on a sed safari to find and replace the new channel names into:

  • The filter ini files
  • various MEDM screens
  • The optic misaligning scripts (which currently live in medm/MISC/ifoalign, and need to get moved to scripts/)
  • A recent BURT snapshot, to restore all of the switches and settings easily. 
  • scripts/activateDQ.py, which is responsible for renaming OL_PIT_IN1_DQ to OPLEV_PERROR, etc.

I've fixed everything that occured to me, and the usual ways I'm used to interacting with the oplevs all seem to work at this time, but it's entirely possible I've overlooked something.

One important note is: because we are now using an effectively immutable QPD library block, the oplev urad conversion has to take place in the DoF matrix. The EPICS records C1:SUS-[OPTIC]_OL_[DOF]_CALIB still exist, but do not multiply the fast signals. Rather, the OL_MTRX elements are multiples of the CALIB value. I thought about making a new QPD_CALIBRATED part or something, but then we're right back to using custom code, which is what we're trying to avoid. 

All of the oplev DoFs are stable, I checked a few loop TFs like ETMY pitch and PRM yaw, and they looked normal. 

  11055   Fri Feb 20 14:44:47 2015 ericqUpdateCDSBad CDS behavior

In an effort to ease the IO load on the framebuilder, I've cleaned up the DQ channels being written to disk. The biggest impact was seven 64kHz channels being written to disk, on ADC channels corresponding to microphones. 

The frame files have gone from 75MB to 57MB. 

  11054   Fri Feb 20 12:29:01 2015 ericqUpdateCDSAll optics damped

I noticed some diagnostic bits in the c1sus IOP complaining about user application timing and FIFO under/overflow (The second and fourth squares next to the DACs on the GDS screen.) Over in crackle-cymac land, I've seen this correlate with large excess DAC noise. After restarting all models, all but one of these is green again, and the optics are now all damped. 

It seems there were some fishy BURT restores, as I found the TT control filters had their inputs and outputs switched off. Some ASS filters were found this way too. More undesired settings may still lurk in the mists...

The interferometer is now realigned, arms locked. 

  11053   Fri Feb 20 12:08:10 2015 ericqUpdateCDSBad CDS behavior

I've been able to get all models running. Most optics are damped, but I'm having trouble with the ITMs, BS and PRM. 

  11052   Thu Feb 19 23:23:52 2015 ChrisUpdateCDSBad CDS behavior

The frontends have some paths NFS-mounted from fb. fb is on the ragged edge of being I/O bound. I'd suggest moving those mounts to chiara. I tried increasing the number of NFS threads on fb (undoing the configuration change I'd previously made here) and it seems to help with EPICS smoothness -- although there are still occasional temporal anomalies in the time channels. The daqd flakiness (which was what led me to throttle NFS on fb in the first place) may now recur as well.

 

Quote:

At about 10AM, the C1LSC frontend stopped reporting any EPICS information. The arms were locked at the time, and remained so for some hours, until I noticed the totally whited-out MEDM screens. The machine would respond to pings, but did not respond to ssh, so we had to manually reboot.

Soon thereafter, we had a global 15min EPICS freeze, and have been in a weird state ever since. Epics has come back (and frozen again), but the fast frontends are still wonky, even when EPICS is not frozen. Intermittantly, the status blinkers and GPS time EPICS values will freeze for multiple seconds at a time, sporadically updating. Looking at a StripTool trace of an IOPs GPS time value shows a line with smooth portions for about 30 seconds, about 2 minutes apart. Between this is totally jagged step function behavior. C1LSC needed to be power cycled again; trying to restart the models is tough, because the EPICS slowdown makes it hard to hit the BURT button, as is needed for the model to start without crashing.

The DAQ network switch, and martian switch inside were power cycled, to little effect. I'm not sure how to diagnose network issues with the frontends. Using iperf, I am able to show hundreds of Mbit/s bandwidth betweem the control room machines and the frontends, but their EPICS is still totally wonky. 

What can we do??? indecision

 

  11051   Thu Feb 19 15:45:43 2015 ericqUpdateCDSBad CDS behavior

At about 10AM, the C1LSC frontend stopped reporting any EPICS information. The arms were locked at the time, and remained so for some hours, until I noticed the totally whited-out MEDM screens. The machine would respond to pings, but did not respond to ssh, so we had to manually reboot.

Soon thereafter, we had a global 15min EPICS freeze, and have been in a weird state ever since. Epics has come back (and frozen again), but the fast frontends are still wonky, even when EPICS is not frozen. Intermittantly, the status blinkers and GPS time EPICS values will freeze for multiple seconds at a time, sporadically updating. Looking at a StripTool trace of an IOPs GPS time value shows a line with smooth portions for about 30 seconds, about 2 minutes apart. Between this is totally jagged step function behavior. C1LSC needed to be power cycled again; trying to restart the models is tough, because the EPICS slowdown makes it hard to hit the BURT button, as is needed for the model to start without crashing.

The DAQ network switch, and martian switch inside were power cycled, to little effect. I'm not sure how to diagnose network issues with the frontends. Using iperf, I am able to show hundreds of Mbit/s bandwidth betweem the control room machines and the frontends, but their EPICS is still totally wonky. 

What can we do??? indecision

  11050   Thu Feb 19 04:16:45 2015 JenneUpdateLSCALS fool attempt with PRFPMI

[Jenne, EricQ]

We tried several times tonight to engage the Fool path with the PRFPMI.  No success. 

First, we locked the arms on ALS, in CARM/DARM mode, and measured the cancellation ability, to make sure that the filter shape and gain was set correctly.  For the PRFPMI, it was okay using the same shape as the single arm case, but the gain was +20.0.  There might be a bit more cancellation to be gained if we adjust the shape at the ~1dB level, but we're already able to get 20dB of cancellation, so we decided that would be enough to give things a try.  To get this much cancellation, we set the phase tracker loops to both have 2kHz UGFs, almost exactly.  We should implement a UGF servo, or the amplitude method version of that as Koji suggested ages ago, so that the phase tracker is always at the same place.

I don't think that REFL 11 is seeing as much CARM as I expect.  We ended up switching over to linearized REFL55 for our attempts.  When we're close to zero CARM offset, the arms are constantly flashing through resonance, and we get the loud buzzing.  REFL11 doesn't seem to see any of this, even though we should be close enough to see some PDH action.  REFL55 does change as we get closer to resonance, so I think it's seeing some real CARM stuff.

We tried engaging the Fool, but I don't think it did anything too useful. We need to make an estimate of what we expect our gain of the REFL loop to be - or at least the sign.

The PRMI is still not stable enough.  It keeps falling out of lock when we get to high-ish arm powers.  Not good.  More brain power tomorrow on the modulation cancellation issue. 

Perhaps if things are stable at moderate arm powers, we can use an excitation to line up the ALS vs. REFL error signals, and then watch the noises of them change as we move around in CARM offset.  This should tell us when the linearized REFL signal is quiet enough that it's worth triggering and trying to transfer over.

 

The last lockloss tonight, there was something funny going on, that we can't explain.  Even though both arms were locked on the CARM/DARM combined ALS signals, beatx doesn't see the giant oscillation that causes carm to lose lock until much later.  Fool was trying to do something, but that should affect both als individal signals in the same way.  Mystery.

Attachment 1: BEATYcrazy_duringCARMdarmLock_18Feb2015.png
BEATYcrazy_duringCARMdarmLock_18Feb2015.png
Attachment 2: 19Feb.png
19Feb.png
  11049   Thu Feb 19 04:09:21 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

X end QPD has recieved 0.2+0.4 absorptive ND filters. Y end QPD got one at 0.6. This appears to have mitigated the saturations for now; the unwhitened signals no longer go negative. The digital gains have been reset. 

 

  11048   Wed Feb 18 19:06:40 2015 KojiUpdateLSCALS Fool impulse response

  11047   Wed Feb 18 17:51:40 2015 ericqUpdateLSCALS Fool impulse response

Koji raised a good question about the step response I wrote about. Namely, if the UGF of the ALS servo is around 100Hz, we would expect it to settle with a characteristic time on the order of tens of milliseconds, not seconds, as was seen in the plot I posted. 

I claim that the reason for the slow response was the fact that the feedforward FM stayed on after the kick, despite the MC filter bank being triggered off. Since it has filters that look like a oscillator at 1Hz, the ringdown or exponential decay of this filter's output in response to the large impulsive output of the PDH control signal just before being triggered down would slowly push the ALS error signal around through the feedforward path. 

Given this reasoning, this should be helped by adding output triggering to the FF filter that uses the MC trigger matrix row, as I wanted to do anyways. I have now added this into the LSC model (as well as DQ at 2kHz for the MC_CTRL_FF_OUT), and the impulse response is indeed much quicker. 

In the following plot, I hit ETMY with a five sample, 5000 amplitude, impulse (rather than a step, as I did yesterday). The system comes back to PDH lock after ~40ms. 

Attachment 1: ALSfool_fasterKick.png
ALSfool_fasterKick.png
  11046   Wed Feb 18 01:58:51 2015 ericqUpdateLSCALS Fool single arm performance

I'm playing around with the lastest ALS fool feedforward on the Yarm, and I like what I'm seeing. 

First, I verified that I could reproduce the TF shapes in ELOG 11041, which I was able to do with a gain of +9.3 and FMs 5 and 6 in the FF module. 

Then, I locked the arm on ALS with full bandwidth, and on POY with the settings currently used the MC module, and took their spectra as references. (I put an excitation on the arm at 443Hz to line them up to the same arbitrary units.)

Then, with ALS at its usual 100Hz UGF and boosts on, the Fool path on, and the MC FM set to trigger on/off at 0.8/0.5, I slowly brought ALS towards zero offset and saw it pop right into resonance. cool I then manually triggered the PDH boosts. 

Here are some spectra, showing that, with the Fool path on:

  • POY unsurprisingly picks up the high frequency noise of the ALS. (Could be mitigated by judicious lowpassing?)
  • The in-fool-loop POY noise is WAY more supressed at low frequencies, so the loops are definitely working together. RMS is about 2x smaller too.

Once the PDH loop is running, the ALS loop can be switched out at the CARM FM output without much of an effect beyond a small kick.


However, looking at the loop shapes, maybe I got lucky here. I took the usual injection TFs at the MC FM, the CARM FM, and at ETMY, to get the overall OLG; all of them have >0.9 coherence pretty much everywhere except the first two points.

 As desired, the PDH loop looks pretty normal.

I have no intuition about how the fooled CARM loop should look, since this is even more complicated than a two-loop system. 

I don't currently know what is causing the odd feature in the overall at ~50Hz, and it spooked me out when I saw the multiple UGF crossings. The only thing I could think of happening there is the pole in the ALS phase tracker boost. I turned it off, and remeasured; the feature persists...


To wrap it up, here's something I think is pretty cool. Here's what happens when I give ETMY a 1000 count position step impulse (no ramp). [Here, CARM is on ALS with G=12, but only FM5 on]

Although the arm was controlled with IR before the kick, ALS maintained control. As soon as ALS brought the arm back towards resonance, the PDH loop picked it right back up.

Coooool.


Some random notes:

  • We should really DQ the output of the feedforward FM. I'll try to remember to do this tonight
  • I was having problems with the zero crossing triggering, so I didn't end up using it, desipte trying to. Maybe we should implement the Schmitt-y style of "Am I below the threshold? Wait. Am I closer to zero now? Ok, Go!"
  • The 1Hz pole in the FF FM rings, unsurprisingly, when the MC FM triggers on briefly, which can be a pain. I made a FM4 which is a Q=1 (instead of 7) pole pair for less ringy. This probably hurts the cancellation somewhat, but I was impatient. 
    • Alternatively, we could try to figure out how to force history clear when the MC filter is triggered off. 

DTT data is attached, in case it's useful to anyone!

Attachment 1: ALSfool_spectra.png
ALSfool_spectra.png
Attachment 2: ALSfool_kick.png
ALSfool_kick.png
Attachment 3: ALSfool_LoopShapes.png
ALSfool_LoopShapes.png
Attachment 4: ALSfool_Feb182015.zip
  11045   Tue Feb 17 19:49:51 2015 KojiUpdateLSCDelay line un-installed again

The modulation setting was reverted.
Demod phase for REFL11/33/55/165 and AS55 were reverted to the previous numbers too.

  11044   Tue Feb 17 16:44:04 2015 KojiUpdateLSCDelay line installed again

For tonight's experiment, I re-installed the delay line cable and changed the attenuation to 10dB for the 55MHz modulation.

I quickly locked the PLL and checked that the modulation is the ratio of the field strength between the worst (19ns) and best
case (28ns) is 31dB, that is ~35 times reduction.

  11043   Tue Feb 17 16:36:08 2015 JenneUpdateLSCALS fool math

Here is an updated cartoon, with the ALS sensor explicitly shown as the beatbox times the closed loop response of the phase tracker servo. 

The most important transfer functions are written on the diagram.  Others can be extracted from the attached Mathematica file (which corresponds to this diagram).

 

Attachment 1: ALSfool_LoopDiagram_17Feb2015.png
ALSfool_LoopDiagram_17Feb2015.png
Attachment 2: ALS_REFL_comboLocking_16Feb2015.nb.zip
  11042   Tue Feb 17 04:04:32 2015 JenneUpdateLSCALS fool math

I re-did the Mathematica notebook according to the most current diagram (note to daytime self: attach .nb file!!!), and found that the denominator has changed, such that plugging in the new D=-A_refl*P_als*S_als gives the same

full-system closed loop gain of    \frac{1}{1-H_{\rm als} - H_{\rm refl} + H_{\rm als}H_{\rm refl}}

where H_{*} = A_* G_* S_* P_*  is the open loop gain, and the * indicates either the REFL or ALS portions of the system. 


I have also plotted some things with Matlab, although I'm a little confused, and my daytime self needs to spend some more time thinking about this.

In the actuators (both for REFL and ALS), I include a pendulum, the digital anti-imaging filters that let us go from the 16kHz model to the 64kHz IOP and the analog anti-imaging filters after the DAC.  Note to self:  still need to include violin filters here.

For the servo gains, I copy the filters that we are using from Foton, and give them the same overall gain multiplier that is in the filter bank.  For the ALS going through the CARM filter bank, this is FMs 1, 2, 3, 5, 6 with a gain of 15.  For the RF (actually, POY here) going through the MC filter bank, this is FMs 4, 5, 7 with a gain of 0.08. 

For the plants of each system, since this is still single arm lock, I just include a cavity pole (80kHz for ALS, 18kHz for REFL). 

In the sensors (both for REFL and ALS), I include the analog anti-aliasing as well as the digital anti-aliasing to allow us to go from the 64kHz IOP to the 16kHz front end system.  For the ALS I also include in the sensor the closed loop response of the phase tracker loop (H/(1-H), where H is the open loop gain of the phase tracker).  For both sensors, I also include a semi-arbitrary number to make the full single-loop open loop gain have a UGF of 200Hz.  In the ALS sensor, I also include a minus sign to make the full open loop gain have the correct phase.

Here I plot the open loop gains of the individual single loops, as well as the open loop gain of the full system (Hals + Hrefl - Hals*Hrefl).  I feel like I must be missing a minus sign in my ALS loop, but I don't know where, and my nighttime brain doesn't want to just throw in minus signs without knowing why.  That will affect how the final ALSfool (blue trace) looks, so maybe it's not really as crazy as it looks right now.

Also, I was trying to explain to myself why we are getting the shape that we are in our measurements of the cancellation (http://nodus.ligo.caltech.edu:8080/40m/11041).  But, I can't.  Below are the plots of the transfer functions from either point 9 or 10 (before or after the G_refl) to point 5, which is the ALS error point.  The measurement in elog 11041 should correspond to the blue trace here.  For these traces, the decoupling is set to just (-A_refl), although there aren't any noticeable changes in the shape if I just set D=0.  If we start with the assumption that D=0, the shape and magnitude are basically identical to this plot, and then as we make D=-A_refl P_als S_als, the transfer functions both go to zero. 

So.  Why is it that with no decoupling, the transfer function from 10 to 5 is tiny?  Why do the shapes plotted below look nothing at all like the measured cancellation shape?  Daytime brain needs to think some more.

Attachment 1: OpenLoopGainComparison_16Feb2015.png
OpenLoopGainComparison_16Feb2015.png
Attachment 2: CancellationTFs_DecouplingIsArefl.png
CancellationTFs_DecouplingIsArefl.png
  11041   Tue Feb 17 00:24:47 2015 rana, jenneUpdateLSCALS Fool filter updated for more cancellation

Today we measured the TFs again and then updated the filter in the POY -> ALS FF path so as to get 10x better cancellation.

The cancellation went from ~10 dB to ~30 dB. This seems good enough. The new filter 'Comp1' is just constructed by eye. We then had to tune the filter module gain to a few %. Seems good enough for now, but we should really try to understand what it is and why it is the way that it is. In the above plot, the ORANGE trace is the old cancellation and the GREEN one is the new one. The filter TF is attached below - its not special, we made it by presing buttons in FOTON until the TF matched the measured TF of ALSY/LSC-MC_CTRL_FF_OUT.

Attachment 1: ALSfoo_150216.png
ALSfoo_150216.png
Attachment 2: 15.png
15.png
  11040   Mon Feb 16 21:52:51 2015 ranaHowToTreasurebig Dataviewer windows

Following this entry, I have made the same change in the controls account on rossa:

In the ~/.grace/gracerc file (create one if it doesn't exist), put in a line which reads:

PAGE LAYOUT FREE

Now we can scale our dataviewer live and playback plots by stretching the window with our mouse. The attached screenshot shows how I filled up one of the vertical monitors with a DV window for arm locking.

Attachment 1: bigDV.png
bigDV.png
  11039   Mon Feb 16 15:08:26 2015 JenneUpdateLSCALS fool measured decoupling TF

Dang it, yes. You're right.  I should have caught that. 

Since Dcpl and Srefl are both zero during the measurement (since it was an ALS lock), this is actually

\frac{{\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als} S_{\rm als}}}{1 - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als} P_{\rm als}}}

So, I need to remove the effect of the ALS closed loop, to get the actual quantity I was looking for.

  11038   Mon Feb 16 03:10:42 2015 KojiUpdateLSCALS fool measured decoupling TF

Wonkey shape: Looks like a loop supression. Your http://nodus.ligo.caltech.edu:8080/40m/11016 also suggests it too, doesn't it?

  11037   Mon Feb 16 02:49:57 2015 JenneUpdateLSCALS fool measured decoupling TF

I have measured very, very carefully the transfer function from pushing on MC2 to the Yarm ALS beatnote.  In the newest loop diagram in http://nodus.ligo.caltech.edu:8080/40m/11030, this is pushing at point 10 and sensing at point 4. 

Since it's a bunch of different transfer functions (to get the high coherence that we need for good cancellation to be possible), I attach my Matlab figure that includes only the useful data points.  I put a coherence cutoff of 0.99, so that (assuming the fit were perfect, which it won't be), we would be able to get a maximum cancellation of a factor of 100. 

This plot also includes the vectfit to the data, which you can see is pretty good, although I need to separately plot the residuals (since the magnitude data is so small, the residuals for the mag don't show up in the auto plot that vectfit gives). 

If you recall from http://nodus.ligo.caltech.edu:8080/40m/11020, we are expecting this transfer function to consist of the suspension actuator (pendulum with complex pole pair around 1Hz), the ALS plant (single pole at 80kHz) and the ALS sensor shape (the phase tracker is an integrator, with a boost consisting of a zero at 666Hz and a pole at 55Hz).  That expected transfer function does not multiply up to give me this wonky shape.  Brain power is needed here.

Just in case you were wondering if this depends on the actuator used (ETM vs MC2), or IFO configuration (single arm vs. PRFPMI), it doesn't.  The only discrepancy between these transfer functions is the expected sign flip between the MC2 and ETMY actuators.  So, they're all pretty consistent. 

Before locking the PRFPMI, I copied the boost filter (666:55) from the YARM ALS over to Xarm ALS, so now both arms have the same boost.

YARM_actTF_compareActuators.pdf


Things to do for ALSfool:

  • Put fitted TF into the MC_CTRL_FF filter bank, and try to measure the expected cancellation, a la http://nodus.ligo.caltech.edu:8080/40m/11009
  • Quick test with single arm, ALS locked using full loop (high gain, all boosts), since the previous versions were with ALS very loosely locked.
    • Does this measured transfer function actually give us good cancellation? 
  • Think.  Why should the transfer function look like this??
  • Try it on the full PRFPMI
Attachment 1: ALSfool_measuredActuatorTF_YarmOnly_15Feb2015.png
ALSfool_measuredActuatorTF_YarmOnly_15Feb2015.png
Attachment 2: YARM_actTF_compareActuators.pdf
YARM_actTF_compareActuators.pdf
  11036   Mon Feb 16 01:45:12 2015 KojiSummaryLSCmodulation depth analysis

Based on the measured modulation profiles, the depth of each modulation was estimated.
Least square sum minimization of the relative error was used for the cost function.
-8th, -12th~-14th, n=>7th are not included in the estimation for the nominal case.
-7th~-9th, -11th~-15th, n=>7th are not included in the estimation for the 3f reduced case.

Nominal modulation

m_f1 = 0.194
m_f2 = 0.234
theta_f1f2 = 41.35deg
m_IMC = 0.00153

3f reduced modulation

m_f1 = 0.191
m_f2 = 0.0579
theta_f1f2 = 180deg
m_IMC = 0.00149

(Sorry! There is no error bars. The data have too few statistics...)

Attachment 1: modulation_nominal.pdf
modulation_nominal.pdf
Attachment 2: modulation_3f_reduced.pdf
modulation_3f_reduced.pdf
  11035   Mon Feb 16 00:08:44 2015 KojiSummaryLSC[ELOG LIST] 3f modulation cancellation

This KTP crystal has the maximum allowed RF power of 10W (=32Vpk) and V_pi = 230V. This corresponds to the maximum allowed
modulation depth of 32*Pi/230 = 0.44. So we probably can achieve gamma_1 of ~0.4 and gamma_2 of ~0.13. That's not x3 but x2,
so sounds not too bad.

Then Kiwamu's triple resonant circuit LIGO-G1000297-v1 actually shows the modulation up to ~0.7. Therefore it is purely an issue
how to deliver sufficient modulation power. (In fact his measurement shows some nonlinearity above the modulation depth of ~0.4
so we should keep the maximum power consumption of 10W at the crystal)

This means that we need to review our RF system (again!)

- Review infamous crazy attn/amp combinations in the frequency generation box.
- Use Teledyne Cougar ampilfier (A2CP2596) right before the triple resonant box. This should be installed closely to the triple resonant box in order to
minimize the effects of the reflection due to imperferct impedance matching.
- Review and refine the triple resonant circuit - it's not built on a PCB but on a universal board. I think that we don't need triple
resonance, but double is OK as the 29.5MHz signal is small.

We want +28V supply at 1X1 for the Teledyne amp and the AOM driver. Do we have any unused Sorensen?

  11034   Sun Feb 15 20:55:48 2015 ranaSummaryLSC[ELOG LIST] 3f modulation cancellation

I wonder if DRMI can be locked on 3f using this lower 55 MHz modulation depth. It seems that PRMI should be unaffected, but that the 3*f2 signals for SRCL will be too puny. Is it really possible to scale up the overall modulation depths by 3x to compensate for this?

  11033   Sun Feb 15 16:20:44 2015 KojiSummaryLSC[ELOG LIST] 3f modulation cancellation

Summary of the ELOGS

3f modulation cancellation theory http://nodus.ligo.caltech.edu:8080/40m/11005

3f modulation cancellation adjustment setup http://nodus.ligo.caltech.edu:8080/40m/11029

Experiment http://nodus.ligo.caltech.edu:8080/40m/11031

Receipe for the 3f modulation cancellation http://nodus.ligo.caltech.edu:8080/40m/11032

Modulation depth analysis http://nodus.ligo.caltech.edu:8080/40m/11036

  11032   Sat Feb 14 22:14:02 2015 KojiSummaryLSC[HOW TO] 3f modulation cancellation

When I finished my measurements, the modulation setup was reverted to the conventional one.
If someone wants to use the 3f cancellation setting, it can be done along with this HOW-TO.


The 3f cancellation can be realized by adding a carefully adjusted delay line and attenuation for the 55MHz modulation
on the frequency generation box at the 1X2 rack.  Here is the procedure:

1) Turn off the frequency generation box

There is a toggle switch at the rear of the unit. It's better to turn it off before any cable action.
The outputs of the frequency generation box are high in general. We don't want to operate
the amplifiers without proper impedance matching in any occasion.

2) Remove the small SMA cable between 55MHz out and 55MHz in (Left arrow in the attachment 1).

According to the photo by Alberto (svn: /docs/upgrade08/RFsystem/frequencyGenerationBox/photos/DSC_2410.JPG),
this 55MHz out is the output of the frequency multiplier. The 55MHz in is the input for the amplifier stages.
Therefore, the cable length between these two connectors changes the relative phase between the modulations at 11MHz and 55MHz.

3) Add a delay line box with cables (Attachment 2).

Connect the cables from the delay line box to the 55MHz in/out connectors. I used 1.5m BNC cables.
The delay line box was set to have 28ns delay.

4) Set the attenuation of the 55MHz EOM drive (Right arrow in the attachment 1) to be 10dB.

Rotate the attenuation for 55MHz EOM from 0dB nominal to 10dB.

5) Turn on the frequency modulation box


For reference, the 3rd attachment shows the characteristics of the delay line cable/box combo when the 3f modualtion reduction
was realized. It had 1.37dB attenuation and +124deg phase shift. This phase change corresponds to the time delay of 48ns.
Note that the response of a short cable used for the measurement has been calibrated out using the CAL function of the network analyzer.

Attachment 1: freq_gen_box.JPG
freq_gen_box.JPG
Attachment 2: delay_line.JPG
delay_line.JPG
Attachment 3: cable_spec.pdf
cable_spec.pdf
  11031   Sat Feb 14 20:37:51 2015 KojiSummaryLSC3f modulation cancellation

Experimental results

- PD response [Attachment 1]

The AUX laser temperature was swept along with the note by Annalisa [http://nodus.ligo.caltech.edu:8080/40m/8369]
It is easier to observe the beat note by closing the PSL shutter as the MC locking yields more fluctuation of the PSL
laser freuqency at low frequency. Once I got the beat note and maximized it, I immediately noticed that the PD response
is not flat. For the next trial, we should use Newfocus 1611. For the measurement today, I decided to characterize the
response by sweeping the beat frequency and use the MAXHOLD function of the spectrum analyzer.

The measured and modelled response of the PD are shown in the attachment 1. It has non-intuitive shape.
Therefore the response is first modelled by two complex pole pair at 127.5MHz with Q of 1, and then the residual was
empirically fitted with 29th polynomial of f.

- Modulation profile of the nominal setting [Attachment 2]

Now the spectrum of the PD output was measured. This is a stiched data of the spectrum between 1~101MHz and 99~199MHz
that was almost simultaneously measured (i.e. Display 1 and Display 2). The IF bandwidth was 1kHz. The PD response correction
described above was applied.

It obviously had the peaks associated with our main modulations. In addition, there are more peaks seen.
The attachment 2 breaks down what is causing the peaks.

  • Carrier: The PLL LO frequency is 95MHz. Therefore the carrier is locked at 95MHz.
  • Modulation sidebands (11/55MHz series):
    Series of sidebands are seen at the both side of the carrier. Their frequency is 95MHz +/- n * fmod  (fmod = 11.066128MHz).
    Note that the sidebands for n>10 were above 200MHz, and n<-9 (indicated in gray) were folded at 0Hz.
    With this measurement BW, the following sidebands were buried in the noise floor.
    n = -8, -12, -13, and -14, n<= -16, and n>=+7
  • Modulation sidebands for IMC and PMC (29.5MHz and 35.5MHz):
    First order sidebands for the IMC and PMC modulations of sidebands are seen at the both side of the carrier.
    Their frequency is 95MHz +/- 29.5MHz or 33.5MHz. The PMC modulation sidebands are supposed to be blocked
    by the PMC. However, due to finite finesse of the PMC, small fraction of the PMC sidebands are transmitted.
    In deed, it is comparable to the modulation depth of the IMC one.
  • RF AM or RF EMI for the main modulation and the IMC modulationand:
    If there is residual RF AM in the PSL beam associated with the IMC and main modulations, it appears as the
    peaks at the modulation frequency and its harmonics. Also EM radiation couples into this measument RF system
    also appears at these frequencies. They are seen at n * fmod  (n=1,2,4,5) and 29.5MHz.
  • Reflection/distortion or leakage from mixer IF to RF:
    The IF port of the mixer naturally has 190MHz signal when the PLL is locked. If the isolation from the IF port to the RF port
    is not enough, this signal can appear in the RF monitor signal via an imperfection of the coupler or a reflection from the PD.
    Also, if the reflecrtion/distortion exist between the PD and the mixer RF input, it also cause the signal around 190MHz.
    It is seen at 190MHz +/- n* fmod. In the plot, the peak at n=0, -1 are visible. In fact these peak were secondarily dominant
    in the spectrum when there was no 6dB attenuation in the PD line. WIth the attenuator, they are well damped and don't disturb
    the main measurment.

From the measured peak height, we are able to estimate the modulation depths for 11MHz, 55MHz, IMC modulations, as well as
the relative phase of the 11MHz and 55MHz modulation. (It is not yet done).

- 3f modulation reduction [Attachment 3]

Now, the redcution of the 3f modulation was tried. The measured modulation levels for the 11MHz and 55MHz were almost the same.
The calculation predicts that the modulation for the 55MHz needs to be 1/3 of the 11MHz one. Therefore the attenuation of 9dB and 10dB
of the modulation attenuation knob at the frequency generation box were tried.

To give the variable delay time in the 55MHz line, EG&G ORTEC delay line unit was used. This allows us to change the delay time from
0ns to 63.5ns with the resolution of 0.5ns. The frequency of 55MHz yields a phase sensitivity of ~20deg/ns (360deg/18ns).
Therefore we can adjust the phase with the precision of 10deg over 1275deg.

The 3rd-order peak at 61.8MHz was observed with measurement span of 1kHz with very narrow BW like 30Hz(? not so sure). The delay
time was swept while measuring the peak height each time. For both the atteuation, the peak height clearly showed the repeatitive dependence
with the period of 18ns, and the 10dB case gave the better result. The difference between the best (1.24e-7 Vpk) and the worst (2.63e-6 Vpk)
was more than a factor of 20.
The 3rd-order peak in the above broadband spectrum measurement was 6.38e-6 Vpk. Considering the attenuation
of the 55MHz modulation by 10dB, we were at the exact unluck phase difference.
The improvement expected from the 3f reduction (in the 33MHz signal)
will be about 50, assuming there is no other coupling mechanism from CARM to REFL33.

I decided to declare the best setting is "10dB attenuation & 28ns delay".

- Resulting modulation profile [Attachment 4]

As a confirmation, the modulation profie was measured as done before the adjustment.
It is clear that the 3rd-order modulation was buried in the floor noise. 10dB attenuation of the 55MHz modulation yields corresponding reduction of the sidebands.
This will impact the signal quality for the 55MHz series error signals, particularly 165MHz ones. We should consider to install the Teledyne Cougar amplifier
next to the EOM so that we can increase the over all modulation depth.

Attachment 1: beat_pd_response.pdf
beat_pd_response.pdf
Attachment 2: beat_nominal.pdf
beat_nominal.pdf
Attachment 3: 3f_reduction.pdf
3f_reduction.pdf
Attachment 4: beat_3f_reduced.pdf
beat_3f_reduced.pdf
  11030   Sat Feb 14 20:20:24 2015 JenneUpdateLSCALS fool cartoon

The ALS fool scheme is now diagrammed up in OmniGraffle, including its new official icon.  The mathematica notebook has not yet been updated.

EDIT, JCD, 17Feb2015:  Updated cartoon and calculation: http://131.215.115.52:8080/40m/11043

 

Attachment 1: ALSfool_LoopDiagram.png
ALSfool_LoopDiagram.png
Attachment 2: ALSfool_LoopDiagram.graffle.zip
  11029   Sat Feb 14 19:54:04 2015 KojiSummaryLSC3f modulation cancellation

Optical Setup

[Attachment 1]

Right before the PSL beam goes into the vacuum chamber, it goes through an AR-wedged plate.
This AR plate produces two beams. One of them is for the IO beam angle/position monitor.
And the other was usually dumped. I decided to use this beam.

A G&H mirror reflects the beam towards the edge of the table.
A 45deg HR mirror brings this beam to the beat set up at the south side of the table.
This beam is S-polarlized as it directly comes from the EOM.

[Attachment 2]

The beam from the PSL goes through a HWP and some matching lenses before the combining beam splitter (50% 45deg P).
The AUX laser beam is attenuated by a HWP and a PBS. The transmitted beam from the PBS is supposed
to have P-polarization. The beam alignment is usually done at the PSL beam side.

The combined beam is steered by a HR mirror and introduced to Thorlabs PDA10CF. As the PD has small diameter
of 0.5mm, the beam needed to be focused by a strong lens.

After careful adjustment of the beam mode matching, polarization, and alignment, the beatnote was ~1Vpp for 2.5Vdc.
In the end, I reduced the AUX laser power such that the beat amplitude went down to ~0.18Vpp (-11dBm at the PD,
-18dBm at the mixer, -27dBm at the spectrum analyzer) in order to minimize nonlinearity of the RF system and
in order that the spectrum analyzer didn't need input attenuation.

Electrical Setup

[Attachment 3]

The PD signal is mixed with a local oscillator signal at 95MHz, and then used to lock the PLL loop.
The PLL loop allows us to observe the peaks with more integration time, and thus with a better signal-to-noise ratio.

The signal from the PD output goes through a DC block, then 6dB attenuator. This attenuator is added to damp reflection
and distortion between the PD and the mixer. When the PLL is locked, the dominant signal is the one at 95MHz. Without this attenuator,
this strong 95MHz signal cause harmonic distortions like 190MHz. As a result, it causes series of spurious peaks at 190MHz +/- n* 11MHz.

10dB coupler is used to peep the PD signal without much disturbing the main line. Considering we have 6dB attanuator,
we can use this coupler output for the PLL and can use the main line for the RF monitor, next time.

The mixer takes the PD signal and the LO signal from Marconi. Marconi is set to have +7dBm output at 95MHz.
FOr the image rejection, SLP1.9 was used. The minicirsuit filters have high-Z at the stop band, we need a 50Ohm temrinator
between the mixer and the LPF.

The error signal from the LPF is fed to SR560 (G=+500, 1Hz 1st-order LPF). I still don't understand why I had to use a LPF
for the locking.
As the NPRO PZT is a frequency actuator, and the PLL is sensitive to the phase, we are supposed to use
a flat response for PLL locking. But it didn't work. Once we check the open loop TF of the system, it will become obvious (but I didn't).

The actuation signal is fed to the fast PZT input of the AUX NPRO laser.
 

Attachment 1: beat_setup1.JPG
beat_setup1.JPG
Attachment 2: beat_setup2.JPG
beat_setup2.JPG
Attachment 3: electrical_setup.pdf
electrical_setup.pdf
  11028   Sat Feb 14 00:48:13 2015 KojiUpdateLSC3f modulation cancellation

[SUCCESS] The 3f sideband cancellation seemed worked nicely.

- Beat effeciency improved: ~30% contrast (no need for amplification)

- PLL locked

- 3f modulation sideband was seen

- The attenuation of the 55MHz modulation and the delay time between the modulation source was adjusted to
have maximum reduction of the 3f sidebands as much as allowed in the setup. This adjustment has been done
at the frequency generation box at 1X2 rack.

- The measurement and receipe for the sideband cancellation come later.


- This means that I jiggled the modulation setup at 1X2 rack. Now the modulation setup was reverted to the original,
but just be careful to any change of the sensing behavior.

- The RF analyzer was returned to the control room.

- The HEPA speed was reduced from 100% (during the action on the table) to 40%.

  11027   Sat Feb 14 00:42:02 2015 KojiUpdateGeneralRF amplifier for ALS

The RF analyzer was returned to the control room. There are two beat notes from X/Y confirmed.

I locked the arms and aligned them with ASS.

When the end greens are locked at TEM00, X/Y beat amplitudes were ~33dBm and ~17dBm. respectively.
I don't judge if they are OK or not, as I don't recall the nominal values.

  11026   Fri Feb 13 19:41:13 2015 ericqUpdateGeneralRF amplifier for ALS

As Koji found one of the spare channels of the ALS/FOL RF amplifier box nonfunctional yesterday, I pulled it out to fix it. I found that one of the sma cables did not conduct.

It was replaced with a new cable from Koji. Also, I rearranged the ports to be consistent across the box, and re-labeled with the gains I observed. 

It has been reinstalled, and the Y frequency counter that is using one of the channels shows a steady beat freq.

I cannot test the amplitude of the green X beat at this time, as Koji is on the PSL table with the PSL shutter closed, and is using the control room spectrum analyzer. However, the dataviewer trace for the fine_phase_out_Hz looks like free swinging cavity motion, so its probably ok.

  11025   Fri Feb 13 18:56:44 2015 ranaUpdateElectronicsSecond QPD Whitening Switch enabled

Depends on the plots of the whitening I guess; if its low freq sat, then we lower the light level with ND filters. If its happening above 10 Hz, then we switch off the whitening.

Quote:

Went to zero CARM offset on ALS; transmission QPDs are still saturating :(

Maybe we need to switch off all whitening.

 

  11024   Fri Feb 13 17:07:51 2015 JenneUpdateElectronicsSecond QPD Whitening Switch enabled

I first updated the DCC branches for the Xend and Yend to reflect the as-built situation from December 2014, and then I updated the drawings after Q's modifications today.

  11023   Fri Feb 13 14:59:13 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

Went to zero CARM offset on ALS; transmission QPDs are still saturating :(

Maybe we need to switch off all whitening.

  11022   Fri Feb 13 14:27:30 2015 ericqUpdateElectronicsSecond QPD Whitening Switch enabled

I have re-enabled the second whitening stage switching on each quadrant of each end's QPD whitening board, to try and avoid saturations at full power. Looking at the spectra while single arm locked, I confirmed that the FM2 whitening switch works as expected. FM1 should be left on, as it is still hard-wired to whiten. 

The oscillations in the Y QPD still exist. Jenne is updating the schematics on the DCC.

  11021   Fri Feb 13 03:44:56 2015 JenneUpdateLSCHeld using ALS for a while at "0" CARM offset with PRMI

[Jenne, Rana]

We wanted to jump right in and see if we were ready to try the new "ALS fool" loop decoupling scheme, so we spent some time with CARM and DARM at "0" offset, held on ALS, with PRMI locked on REFL33I&Q (no offsets).  Spoiler alert:  we weren't ready for the jump.

The REFL11 and AS55 PDs had 0dB analog whitening, which means that we weren't well-matching our noise levels between the PD noise and the ADC noise.  The photodiodes have something of the order nanovolt level noise, while the ADC has something of the order microvolt level noise.  So, we expect to need an analog gain of 1000 somewhere, to make these match up.  Anyhow, we have set both REFL11 and AS55 to 18dB gain. 

On a related note, it seems not so great for the POX and POY ADC channels to be constantly saturated when we have some recycling gain, so we turned their analog gains down from 45dB to 0dB.  After we finished with full IFO locking, they were returned to their nominal 45dB levels. 

We also checked the REFL33 demod phase at a variety of CARM offsets, and we see that perhaps it changes by one or two degrees for optimal rotation, but it's not changing drastically.  So, we can set the REFL33 demod phase at large CARM offset, and trust it at small CARM offset.

We then had a look at the transmon QPD inputs (before the dewhitening) for each quadrant.  They are super-duper saturating, which is not so excellent. 
QPDsaturation_12Feb2015.pdf

We think that we want to undo the permanently-on whitening situation.  We want to make the second stage of whitening back to being switchable.  This means taking out the little u-shaped wires that are pulling the logic input of the switches to ground.  We think that we should be okay with one always on, and one switchable.  After the modification, we must check to make sure that the switching behaves as expected.  Also, I need to figure out what the current situation is for the end QPDs, and make sure that the DCC document tree matches reality.  In particular, the Yend DCC leaf doesn't include the gain changes, and the Xend leaf which does show those changes has the wrong value for the gain resistor.

After this, we started re-looking at the single arm cancellation, as Rana elogged about separately.

ALSfool_12Feb2015.pdf

Attachment 1: QPDsaturation_12Feb2015.pdf
QPDsaturation_12Feb2015.pdf
Attachment 2: ALSfool_12Feb2015.pdf
ALSfool_12Feb2015.pdf
  11020   Fri Feb 13 03:28:34 2015 ranaUpdateLSCNew Locking Paradigm - Loop-gebra

Not so fast!

In the drawing, the FF path should actually be summed in after the Phase Tracker (i.e. after S_ALS). This means that the slow response of the phase tracker needs to be taken into account in the FF cancellation filter. i.e. D = -A_REFL * P_ALS * S_ALS. Since the Phase Tracker is a 1/f loop with a 1 kHz UGF, at 100 Hz, we can only get a cancellation factor of ~10.

So, tonight we added a 666:55 boost filter into the phase tracker filter bank. I think this might even make the ALS locking loops less laggy. The boost is made to give us better tracking below ~200 Hz where we want better phase performance in the ALS and more cancellation of the ALS-Fool. If it seems to work out well we can keep it. If it makes ALS more buggy, we can just shut it off.

Its time to take this loop cartoon into OmniGraffle.

  11019   Thu Feb 12 23:47:45 2015 KojiUpdateLSC3f modulation cancellation

- I built another beat setup on the PSL table at the South East side of the table.
- The main beam is not touched, no RF signal is touched, but recognize that I was present at the PSL table.
- The beat note is found. The 3rd order sideband was not seen so far.
- A PLL will be built tomorrow. The amplifier box Manasa made will be inspected tomorrow.

- One of the two beams from the picked-off beam from the main beam line was introduced to the beat setup.
(The other beam is used of for the beam pointing monitors)
There is another laser at that corner and the output from this beam is introduced into the beat setup.
The combined beam is introduced to PDA10CF (~150MHz BW).

- The matching of the beam there is poor. But without much effort I found the beat note.
  The PSL laser had 31.33 deg Xtal temp. When the beat was found, the aux laser had the Xtal temp of 40.88.

- I could observe the sidebands easily, with a narrower BW of the RF analizer I could see the sidebands up to the 2nd order.
  The 3rd order was not seen at all.

- The beat note had the amplitude of about -30dBm. One possibility is to amplify the signal. I wanted to use a spare channel
of the ALS/FOLL amplifier box. But it gave me rather attenuation than any amplification.
I'll look at the box tomorrow.

- Also the matching of two beams are not great. The PD also has clipping I guess. These will also be improved tomorrow

- Then the beat note will be locked at a certain frequency using PLL so that we can reduce the measurement BW more.

 

  11018   Thu Feb 12 23:40:12 2015 ranaUpdateIOOBounce / Roll stopband filters added to MC ASC filter banks

The filters were already in the damping loops but missing the MC WFS path. I checked that these accurately cover the peaks at 16.5 Hz and 23.90 and 24.06 Hz.

Attachment 1: 59.png
59.png
ELOG V3.1.3-